if x^10 * x^2a= x^100, what is the value of a?
Found 2 solutions by stanbon, josmiceli:
Answer by stanbon(75887)

if x^10 * x^2a= x^100, what is the value of a?
-------------
x^(10+2a) = x^100
---
10+2a = 100
2a = 90
a = 45
===========
Cheers,
stan H.
===========

Answer by josmiceli(19441)